OpenStudy (loser66):

Let f: G--> H be a surjective homomorphism of groups. Prove that H is abelian iff the commutator subgroup G' is contained in the kernel of f. Please, help

OpenStudy (zarkon):

where are you stuck?

OpenStudy (loser66):

I am stuck at commutator subgroup

OpenStudy (zarkon):

ok...what about the commutator subgroup are you stuck on?

OpenStudy (loser66):

by definition, G'={aba^-b^- | a, b in G}, if it is in kernel of f, then f (G') = e_H

OpenStudy (loser66):

and f(G') = f(aba^-b^-) = f(a) f(b) f(a^-) f(b^-) = e_H I don't see how to force H is abelian from this

OpenStudy (zarkon):

\[f(a^{-1})=f(a)^{-1}\]

OpenStudy (zarkon):

also f is surjective

OpenStudy (loser66):

yes, but if H is abelian, then we can change their location to get e_H. We don't have it yet.

OpenStudy (loser66):

all we can do is f(ab) f(ba)^- = e_H

OpenStudy (zarkon):

which direction are you trying to prove right now

OpenStudy (loser66):

I 've just start, either way is ok. Now, I do if commutator subgroup G' in the kernel, then H is abelian.

OpenStudy (zarkon):

ok so we have \[f(aba^{-1}b^{-1})=ker(f)\] right

OpenStudy (zarkon):

not equal

OpenStudy (zarkon):

element of

OpenStudy (loser66):

yes

OpenStudy (zarkon):

\[f(aba^{-1}b^{-1})\in ker(f)\]

OpenStudy (loser66):

yes,

OpenStudy (loser66):

oh, no

OpenStudy (zarkon):

lol

OpenStudy (loser66):

aba^- b^- in ker f, not f(aba^-b^-) in kernel f

OpenStudy (zarkon):

aba^-1b^-1 is is in the ker

OpenStudy (loser66):

yes, :)

OpenStudy (zarkon):

either way... \[f(aba^{-1}b^{-1})=e_H\] ok :)

OpenStudy (loser66):

yes,

OpenStudy (zarkon):

now break it up as \[f(a)f(b)f(a^{-1})f(b^{-1})=e_{H}\]

OpenStudy (zarkon):

ok?

OpenStudy (loser66):

yes

OpenStudy (zarkon):

then \[f(a)f(b)f(a)^{-1}f(b)^{-1}=e_{H}\]

OpenStudy (zarkon):

ok?

OpenStudy (loser66):

ok

OpenStudy (zarkon):

then \[f(a)f(b)=f(b)^{-1}f(a)^{-1}\]

OpenStudy (loser66):

yes

OpenStudy (zarkon):

that last step was a typo (copied too much) \[f(a)f(b)f(a)^{-1}f(b)^{-1}=e_{H}\] implies \[f(a)f(b)=f(b)f(a)\]

OpenStudy (zarkon):

\[f(a)f(b)f(a)^{-1}f(b)^{-1}=e_{H}\] \[f(a)f(b)f(a)^{-1}f(b)^{-1}f(b)f(a)=e_{H}f(b)f(a)\]

OpenStudy (loser66):

yes, I see it

OpenStudy (loser66):

that shows H is abelian, lalala, right?

OpenStudy (zarkon):

yes because f is surjective

OpenStudy (loser66):

how it relate to surjective?

OpenStudy (zarkon):

every two elements of H can be written as f(a) or f(b)

OpenStudy (zarkon):

for any \[\alpha,\beta\in H\] there exists \[a,b\in G\] such that \[f(a)=\alpha\] and \[f(b)=\beta\]

OpenStudy (zarkon):

and that is true because f is surjective

OpenStudy (loser66):

backward, :) for any f(a) , f(b) in H, there exists a, b in G. That's the definition of surjective, right?

OpenStudy (loser66):

lol, yes, my bad,

OpenStudy (loser66):

I see it

OpenStudy (zarkon):

that is what I said...for any elements in H there are elements in G that are mapped to the elements in H

OpenStudy (loser66):

The converse part is backward of what we did, right? (no, you did, not me , lol)

OpenStudy (zarkon):

pretty much

OpenStudy (loser66):

Thanks a lot. I am so surprised that I can get your help. Honestly, I never think I can get this offer. lol.

OpenStudy (zarkon):

why is that?

OpenStudy (loser66):

How to say!! to me, you wander and rarely comment on the post.

OpenStudy (zarkon):

I have to be in the mood. I do math all day. Sometimes I just like to watch others do it

OpenStudy (loser66):

and you are not allowed other send message to get your help. To me, it 's a signal of "let me alone" hahaha

OpenStudy (zarkon):

A large percentage of the time I have a solution to the persons problem but I just want to see what other people come up with

OpenStudy (loser66):

Sometimes, I got mad at .... you, hihihi.... since you are one of the very few people in this site know the hardest stuff, but you don't help.

OpenStudy (zarkon):

lol ;)

OpenStudy (loser66):

Whatever, I won lottery ticket today. hahaha... that price makes my mood up.

OpenStudy (zarkon):

:)

OpenStudy (loser66):

Again, thanks a lot, my lottery ticket.

OpenStudy (zarkon):

no problem ... later